# Quandles as pre-Lie skew braces, set-theoretic Hopf algebras &\&& universal ${\cal R}$-matrices

### v1

@@ 5.1. The quandle Hopf algebra
Let $(X,\triangleright)$ be a quandle and $(X,\bullet)$ be a group, such that $a\tiny\bullet b=b\bullet(b\triangleright a).$ Let also ${\cal A}$ be the quandle algebra (Definition 5.1), ${\cal R}=\sum_{a}h_{a}\otimes q_{a}$ is a solution of the Yang-Baxter equation and $q_{a}:X\to{\cal A}$ is a strong group homomorphism (i.e. $q_{a}q_{b}=q_{a\bullet b}$)
Then the structure $({\cal A},\Delta,\epsilon,S,{\cal R})$ is a quasi-triangular Hopf algebra if $h_{a}^{2}=h_{a},$ for all $a\in X$:
  • Co-product. $\Delta(q_{a})=q_{a}\otimes q_{a},\quad\Delta(q^{-1}_{a})=q^{-1}_{a}\otimes q^{-1}_{a},\quad\Delta(h_{a})=\sum_{b,c\in X}h_{b}\otimes h_{c}\Big{|}_{b\bullet c=a}.$
  • Co-unit. $\epsilon(q_{a})=1,$ $\epsilon(h_{a})=\delta_{a,1}.$
  • Antipode. $S(q_{a})=q_{a}^{-1},$ $S(h_{a})=h_{a^{*}},$ where $a^{*}$ is the inverse in $(X,\bullet)$ for all $a\in X.$
@@
We recall that the maps $\Delta,\epsilon$ and $S$ are algebra homomorphisms, consistent with the relations of the quandle algebra. We are now going to show all the axioms of a quasi-triangular Hopf algebra. We first identify,
${\cal R}_{13}{\cal R}_{12}=\sum_{a\in X}h_{a}^{2}\otimes q_{a}\otimes q_{a}=:\sum_{a\in X}h_{a}^{2}\otimes\Delta(q_{a})=(\mbox{id}\otimes\Delta){\cal R},\quad\mbox{if}\quad ha^{2}=h_{a}$ (5.10)
@@
hence we immediately read of $\Delta(h_{a}),\ \Delta(q_{a})$ as ($\Delta:{\cal A}\to{\cal A}\otimes{\cal A}$ is by construction an algebra homomorphism, but this can be also explicitly checked):
@@
where $\Delta^{(op)}=\sigma\circ\Delta,$ $\sigma$ is the flip map.
@@
    $\sum_{a,b\in X}\epsilon(h_{a})h_{b}=\sum_{a,b}h_{a}\epsilon(h_{b})\Big{|}_{a\bullet b=c}=h_{c}\Rightarrow\epsilon(h_{a})=\delta_{a,1}.$ (5.13)
@@
    For $q_{a},$ we immediately have $S(q_{a})=q_{a}^{-1}$ and (recall $h_{a}h_{b}=\delta_{a,b}h^{2}_{a}$ and $\sum_{a\in X}h^{2}_{a}=1_{\cal A}$)
    $\sum_{a,b\in X}S(h_{a})h_{b}\Big{|}_{a\bullet b=c}=\sum_{a,b\in X}h_{a}S(h_{b})\Big{|}_{a\bullet b=c}=\delta_{c,1}1_{\cal A}\ \Rightarrow\ S(h_{a})=h_{a^{*}},\$
@@
    We conclude that $({\cal A},\Delta,\epsilon,S)$ is a Hopf algebra, but $({\cal A},\Delta,\epsilon,S,{\cal R})$ is not a quasi-triangular Hopf algebra if $h_{a}^{2}\neq h_{a}$ due to equ. (5.10), that is all the axioms of a quasi-triangular Hopf algebra are satisfied except ${\cal R}_{13}{\cal R}_{12}=(\mbox{id}\otimes\Delta){\cal R}.$ If $h_{a}^{2}=h_{a}$ then a quasi-triangular Hopf algebra is recovered. Henceforth, we consider $h_{a}^{2}=h_{a},$ for all $a\in X.$
@@
Let $(X,\triangleright)$ be a quandle and $(X,\bullet)$ be a magma such that $a\tiny\bullet b=b\bullet(b\triangleright a).$ Let also ${\cal A}$ be the quandle algebra (Definition 5.1), ${\cal R}=\sum_{a}h_{a}\otimes q_{a}$ is a solution of the Yang-Baxter equation and $q_{a}:X\to{\cal A}$ is a weak algebra homomorphism (Definition 5.4). Then the structure $({\cal A},\Delta,\epsilon,{\cal R})$ is a quasi-triangular quasi-bialgebra.
@@
[…]
Hence we immediately read of $\Delta(h_{a}),\ \Delta(q_{a})$ as ($\Delta:{\cal A}\to{\cal A}\otimes{\cal A}$ is by construction an algebra homomorphism, but this can be also explicitly checked):
@@
From the Yang-Baxter equation we have $\leavevmode\nobreak\ \Delta^{(op)}(q_{a}){\cal R}=\Delta(q_{a}){\cal R},$   $\Delta^{(op)}(h_{a}){\cal R}={\cal R}\Delta(h_{a}),$ where $\Delta^{(op)}=\sigma\circ\Delta,$ $\sigma$ is the flip map.
@@
[…]
    Notice that expressions (5.17) and (5.18) are in general distinct, given that the operation $\bullet$ is not necessarily associative.
@@
[…]
    The generators $q_{a}$ are group-like elements, so $\epsilon(q_{a})=1.$ For the generators $h_{a},$ the following must be satisfied:
    $\sum_{a,b\in X}\epsilon(h_{a})h_{b}=\sum_{a,b}h_{a}\epsilon(h_{b})\Big{|}_{a\bullet b=c}=h_{c}.$ (5.19)
    Note however that if a left neutral element $e$ exists and if $\epsilon(h_{a})=\delta_{a,e},$ then
[…]
    which suggests that we are dealing with a quasi-bialgebra.
Moreover, because of (5.15) and the lack of coassociativity of $\Delta(h_{a})$ we conclude that $({\cal A},\Delta,\epsilon)$ is not a bialgebra, but rather a quasi-bialgebra, which needs to be fully identified by deriving the coassociator (see also relevant study for the set-theoretic quasi-bialgebra [16, 17]). This analysis however will be presented in detail in a future work.
5.2. The set-theoretic Hopf algebra
In this subsection we suitably extend the quandle Hopf algebra ${\cal A}$ in order to construct the quasi-triangular (quasi) Hopf algebra associated to general set-theoretic solutions of the Yang-Baxter equation.
We start our analysis by defining the set-theoretic Yang-Baxter algebra (see also [3]).
@@ 5.2. The set-theoretic Hopf algebra
(The set-theoretic Yang-Baxter algebra.) Let ${\cal A}$ be the quandle algebra (Definition 5.1). Let also $(X,\circ,1)$ be a group, such that for all $a,b\in X,$ $a\circ b=\sigma_{a}(b)\circ\tau_{b}(a),$ where $\sigma_{a},\ \tau_{b}:X\to X,$ and $\sigma_{a}$ is a bijection. We say that the unital, associative algebra $\hat{\cal A}$ over $k,$ generated by indeterminates $1_{\hat{\cal A}},q_{a},q_{a}^{-1},h_{a},\in{\cal A}$ ($h_{a}=h_{b}\Leftrightarrow a=b$) and $w_{a},w^{-1}_{a}\in\hat{\cal A},$ $a\in X,$ and relations, $a,b\in X:$
$q_{a}q_{a}^{-1}=1_{\hat{\cal A}},\quad q_{a}q_{b}=q_{b}q_{b\triangleright a},\quad h_{a}h_{b}=\delta_{a,b}h_{a},\quad q_{b}h_{b\triangleright a}=h_{a}q_{b},$
$w_{a}w_{a}^{-1}=1_{\hat{\cal A}},\quad w_{a}w_{b}=w_{\sigma_{a}(b)}w_{\tau_{b}(a)}\quad w_{a}h_{b}=h_{\sigma_{a}(b)}w_{a},\quad w_{a}q_{b}=q_{\sigma_{a}(b)}w_{a}$ (5.21)
is a set-theoretic Yang-Baxter algebra.
@@
Let $c=\sum_{a\in X}h_{a},$ then $c$ is a central element of the set-theoretic Yang-Baxter algebra $\hat{\cal A}$.
@@
Let $\hat{\cal A}$ be the set-theoretic Yang-Baxter algebra (Definition 5.13), then for all $a,b,c\in X,$
@@
Let $\hat{\cal A}$ be the set-theoretic Yang-Baxter algebra, then $q,w:X\to\hat{\cal A},$ $a\mapsto q_{a},\ w_{a}$ are called:
  1. (1)
    Strong $(X,\bullet)-$homomorphism and strong $(X,\circ)-$homomorphism respectively if $q_{a}q_{b}=q_{a\bullet b},$ and $w_{a}w_{b}=w_{a\circ b},$ for all $a,b\in X$.
  2. (2)
    Weak $(X,\bullet)-$homomorphism and weak $(X,\circ)-$homomorphism respectively if there exist maps $F,\ G:X\to{\cal A},$ $a\mapsto F_{a},\ G_{a},$ such that $q_{a}q_{b}=F_{a\bullet b}$ and $w_{a}w_{b}=G_{a\circ b},$ for all $a,b\in X$. ($F\neq q$, $G\neq w$; if $F=Q,\ G=w,$ then one recovers a strong homomorphism).
@@
(Hopf algebra) Let $\hat{\cal A}$ be the set-theoretic Yang-Baxter algebra (Definition 5.13), ${\cal R}=\sum_{b}h_{b}\otimes q_{b}$ be the quandle universal ${\cal R}$-matrix, and $({\cal A},\Delta,\epsilon,S,{\cal R})$ is the quasi-triangular Hopf algebra of Theorem 5.11. Let also for all $a,b,x\in X,$ $\sigma_{1}(b)=b$ and
@@
Proposition 5.17 can be generalized for $(X,\bullet)$ being a magma (recall $(X,\circ,1)$ is a group: $a\circ b=\sigma_{a}(b)\circ\tau_{b}(a)$), if for all $x\in X,$ $\exists g_{x}:X\to X,$ $a\mapsto g_{x}(a),$ being a bijection, such that for all $a,b,x\in X,$ $\sigma_{1}(b)=b=g_{1}(b)$ and
@@
Note that in this case according to Remark 5.12 $(\hat{\cal A},\Delta,\epsilon)$ is a quasi-bialgebra.
We now introduce some handy notation that can be used in the following: let $i,j,k\in\{1,2,3\},$ then ${\cal F}_{jik}=\sigma_{ij}\circ{\cal F}_{ijk}$ and ${\cal F}_{ikj}=\sigma_{jk}\circ{\cal F}_{ijk},$ where $\sigma$ is the flip map.
@@
(Drinfel’d twist [21]) Let ${\cal R}=\sum_{a\in X}h_{a}\otimes q_{a}\in{\cal A}\otimes{\cal A}$ be the quandle universal ${\cal R}-$matrix, ${\cal F}\in\hat{\cal A}\otimes\hat{\cal A},$ such that ${\cal F}=\sum_{b\in X}h_{b}\otimes w_{b}^{-1},$ ${\cal R}_{ij}^{F}:={\cal F}_{ji}{\cal R}_{ij}{\cal F}_{ij}^{-1},$ $i,j\in\{1,2,3\}$ and $w:X\to\hat{\cal A}$ is a weak $(X,\circ)-$homomorphism. We also define:
@@
    Indeed, this is proved by a direct computation and use of the set-theoretic Yang-Baxter algebra. In fact, ${\cal F}_{123}=\sum_{a,b\in X}h_{a}\otimes h_{b}w_{a}^{-1}\otimes w_{b}^{-1}w_{a}^{-1}.$
@@
    (ii) By means of the relations of the algebra $\hat{\cal A}$ we compute, using also that $w_{a}w_{b}=F_{a\circ b},$ for some map $F:X\to\hat{\cal A}.$
    ${\cal F}^{*}_{21,3}{\cal R}_{12}=\sum_{a,c\in X}h_{a}\otimes q_{a}h_{a\triangleright c}\otimes F^{-1}_{c\circ\sigma^{-1}_{c}(a)},\quad{\cal R}_{12}{\cal F}^{*}_{12,3}=\sum_{a,b\in X}h_{a}\otimes q_{a}h_{\sigma_{a}(b)}\otimes F^{-1}_{a\circ b}.$
    Due to the fact that $\leavevmode\nobreak\ b\triangleright a=\sigma_{b}(\tau_{\sigma^{-1}_{a}(b)}(a))$ and $a\circ b=\sigma_{a}(b)\circ\tau_{b}(a)$ we conclude that ${\cal F}^{*}_{21,3}{\cal R}_{12}={\cal R}_{12}{\cal F}^{*}_{12,3}$ and consequently (recall $F_{213}={\cal F}_{21}{\cal F}^{*}_{21,3}$)
    ${\cal F}_{213}{\cal R}_{12}={\cal F}_{21}{\cal F}^{*}_{21,3}{\cal R}_{12}={\cal F}_{21}{\cal R}_{12}{\cal F}^{*}_{12,3}={\cal R}_{12}^{F}{\cal F}_{123}.$
@@
[21] Let ${\cal F}$ be an admissible twist and ${\cal R}$ be a solution of the Yang-Baxter equation. Then ${\cal R}^{F}:={\cal F}^{(op)}{\cal R}{\cal F}^{-1}$ (${\cal F}^{(op)}=\sigma\circ{\cal F}$) is also a solution of the Yang-Baxter equation.
@@
We are now going to examine the twisted ${\cal R}-$matrix as well as the twisted co-products (see also [16, 17, 18]).
@@
(Twisted universal ${\cal R}$-matrix) It is worth extracting the explicit expressions of the twisted universal ${\cal R}-$matrix and the twisted coproducts of the algebra. We recall the admissible twist ${\cal F}=\sum_{b\in X}h_{b}\otimes w_{b}^{-1}.$ We consider the more general case where $q$ is a weak $(X,\bullet)-$homomorphism (Remark 5.12).
@@
    Then, the twisted coproducts read as
    $\Delta_{F}(w_{a})=\sum_{b\in X}h_{\sigma_{a}(b)}w_{a}\otimes w_{\tau_{b}(a)},$
    $\Delta_{F}(h_{a})=\sum_{b\in X}h_{b}\otimes w_{b}^{-1}h_{c}w_{b}\Big{|}_{b\bullet c=a},$
    $\Delta_{F}(q_{a})=\sum_{b\in X}q_{a}h_{a\triangleright b}\otimes w_{b}^{-1}q_{a}w_{a\triangleright b},$

### v3

@@ 5.1. The rack & quandle algebras
Let ${\cal A}$ be the quandle algebra (Definition 5.2), with $(X,\bullet,e)$ being a group. Let also ${\cal R}=\sum_{a\in X}h_{a}\otimes q_{a}$ be a solution of the Yang-Baxter equation and $q_{a}:X\to{\cal A}$ be a strong group homomorphism (i.e. $q_{a}q_{b}=q_{a\bullet b}$).
Then the structure $({\cal A},\Delta,\epsilon,S,{\cal R})$ is a quasi-triangular Hopf algebra:
  • Co-product. $\Delta:{\cal A}\to{\cal A}\otimes{\cal A},$ $\leavevmode\nobreak\ \Delta(q_{a}^{\pm 1})=q_{a}^{\pm 1}\otimes q_{a}^{\pm 1}$ and $\Delta(h_{a})=\sum_{b,c\in X}h_{b}\otimes h_{c}\Big{|}_{b\bullet c=a}.$
  • Co-unit. $\leavevmode\nobreak\ \epsilon:{\cal A}\to k,$ $\epsilon(q_{a}^{\pm 1})=1,$ $\leavevmode\nobreak\ \epsilon(h_{a})=\delta_{a,e}.$
  • Antipode. $\leavevmode\nobreak\ S:{\cal A}\to{\cal A},$ $\leavevmode\nobreak\ S(q_{a}^{\pm 1})=q_{a}^{\mp 1},$ $S(h_{a})=h_{a^{*}},$ where $a^{*}$ is the inverse in $(X,\bullet)$ for all $a\in X.$
@@
We recall that the maps $\Delta,\epsilon$ are algebra homomorphisms and $S$ is an anti-homomorphism, consistent with the relations of the quandle algebra. We are now going to show all the axioms of a quasi-triangular Hopf algebra. We first identify,
${\cal R}_{13}{\cal R}_{12}=\sum_{a\in X}h_{a}\otimes q_{a}\otimes q_{a}=:\sum_{a\in X}h_{a}\otimes\Delta(q_{a})=(\mbox{id}\otimes\Delta){\cal R},$ (5.10)
@@
hence we immediately read of $\Delta(h_{a}),\ \Delta(q_{a})$ as ($\Delta:{\cal A}\to{\cal A}\otimes{\cal A}$ should be an algebra homomorphism, but this can be explicitly checked via the distributivity condition $a\triangleright(b\bullet c)=(a\triangleright b)\bullet(a\triangleright c),$ which readily follows from, $a\triangleright b=a^{*}\bullet b\bullet a$):
@@
where $\Delta^{(op)}=\pi\circ\Delta,$ $\pi$ is the flip map.
@@
    $\sum_{a,b\in X}\epsilon(h_{a})h_{b}=\sum_{a,b}h_{a}\epsilon(h_{b})\Big{|}_{a\bullet b=c}=h_{c}\Rightarrow\epsilon(h_{a})=\delta_{a,e}.$ (5.13)
@@
    For $q_{a},$ we immediately have $S(q_{a})=q_{a}^{-1}$ and (recall $h_{a}h_{b}=\delta_{a,b}h_{a}$ and $\sum_{a\in X}h_{a}=1_{\cal A}$)
    $\sum_{a,b\in X}S(h_{a})h_{b}\Big{|}_{a\bullet b=c}=\sum_{a,b\in X}h_{a}S(h_{b})\Big{|}_{a\bullet b=c}=\delta_{c,e}1_{\cal A}\ \Rightarrow\ S(h_{a})=h_{a^{*}},\$
@@
    We conclude that $({\cal A},\Delta,\epsilon,S,{\cal R})$ is a quasi-triangular Hopf algebra. ∎
Corollary 5.13.
Let ${\cal A}$ be the rack algebra, then the subalgebra consisting of the indeterminates $1_{\cal A},\ q_{a}\ q_{a}^{-1},$ and relations for all $a\in X$
$q_{a}^{-1}q_{a}=q_{a}q_{a}^{-1}=1_{\cal A},\quad q_{a}q_{b}=q_{b}q_{b\triangleright a},$ (5.14)
is a Hopf algebra with:
  1. (1)
    Co-product. $\Delta:{\cal A}\to{\cal A}\otimes{\cal A},$ $\leavevmode\nobreak\ \Delta(q_{a}^{\pm 1})=q_{a}^{\pm 1}\otimes q_{a}^{\pm 1}.$
  2. (2)
    Co-unit. $\epsilon:{\cal A}\to k,$ $\leavevmode\nobreak\ \epsilon(q_{a}^{\pm 1})=1.$
  3. (3)
    Antipode. $S:{\cal A}\to{\cal A},$ $\leavevmode\nobreak\ S(q_{a}^{\pm 1})=q_{a}^{\mp 1}.$
Proof.
The proof is straightforward by Theorem 5.12. ∎
@@
Let $(X,\triangleright)$ be a quandle and $(X,\bullet)$ be a magma with a left neutral element, such that $a\bullet b=b\bullet(b\triangleright a).$ Let also ${\cal A}$ be the quandle algebra (Definition 5.2), ${\cal R}=\sum_{a\in X}h_{a}\otimes q_{a}$ is a solution of the Yang-Baxter equation and $q_{a}:X\to{\cal A}$ is a weak algebra homomorphism (Definition 5.5) and $a\triangleright(b\bullet c)=(a\triangleright b)\bullet(a\triangleright c).$ Then the structure $({\cal A},\Delta,\epsilon,{\cal R})$ is a quasi-triangular quasi-bialgebra.
@@
${\cal R}_{13}{\cal R}_{12}=\sum_{a\in X}h_{a}\otimes q_{a}\otimes q_{a}=:\sum_{a}h_{a}\otimes\Delta(q_{a})=(\mbox{id}\otimes\Delta){\cal R},$ (5.15)
${\cal R}_{13}{\cal R}_{23}=\sum_{a,b\in X}h_{a}\otimes h_{b}\otimes F_{c}\Big{|}_{a\bullet b=c}=:\sum_{c\in X}\Delta(h_{c})\otimes F_{c}\neq(\Delta\otimes\mbox{id}){\cal R},\leavevmode\nobreak\ (F\neq q).$ (5.16)
Hence we immediately read of $\Delta(h_{a}),\ \Delta(q_{a})$ as:
@@
From the Yang-Baxter equation and (5.15), (5.16) we conclude $\leavevmode\nobreak\ \Delta^{(op)}(q_{a}){\cal R}=\Delta(q_{a}){\cal R},$   $\Delta^{(op)}(h_{a}){\cal R}={\cal R}\Delta(h_{a}),$ where $\Delta^{(op)}=\pi\circ\Delta,$ $\pi$ is the flip map.
@@
    Co-associativity.: $(\mbox{id}\otimes\Delta)\Delta=(\Delta\otimes\mbox{id})\Delta.$ Coassociativity for $\Delta(q_{a}^{\pm 1})$ is obvious, but
    $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ (\Delta\otimes\mbox{id})\Delta(h_{a})=\sum_{b,c,d\in X}h_{b}\otimes h_{c}\otimes h_{d}\Big{|}_{(b\bullet c)\bullet d=a},\leavevmode\nobreak\ \leavevmode\nobreak\ (\mbox{id}\otimes\Delta)\Delta(h_{a})=\sum_{b,c,d\in X}h_{b}\otimes h_{c}\otimes h_{d}\Big{|}_{b\bullet(c\bullet d)=a}.$
    Notice that expressions above are in general distinct, given that the operation $\bullet$ is not necessarily associative.
@@
    The generators $q_{a}$ are group-like elements, so $\epsilon(q_{a})=1.$ For the generators $h_{a},$ the following must be computed: $\sum_{a,b\in X}\epsilon(h_{a})h_{b}\Big{|}_{a\bullet b=c}$ and $\quad\sum_{a,b}h_{a}\epsilon(h_{b})\Big{|}_{a\bullet b=c}.$ Note that if a left neutral element $e$ exists and if $\epsilon(h_{a})=\delta_{a,e},$ then $(\epsilon\otimes\mbox{id})\Delta(h_{c})=\sum_{a,b\in X}\epsilon(h_{a})h_{b}\Big{|}_{a\bullet b=c}=h_{c},$ which suggests that in this case we are dealing with a quasi-bialgebra.
Moreover, because of (5.15), (5.16) and the lack of coassociativity of $\Delta(h_{a})$ we conclude that $({\cal A},\Delta,\epsilon)$ is not a bialgebra, but rather a quasi-bialgebra, which needs to be fully identified by deriving the coassociator (see also relevant study for the set-theoretic quasi-bialgebra [21, 22]). This analysis however will be presented in detail in a future work.
5.2. The set-theoretic algebras
In this subsection we suitably extend the rack and quandle algebras in order to construct the universal ${\cal R}-$matrix associated to general set-theoretic solutions of the Yang-Baxter equation.
We first define the decorated rack algebra and the set-theoretic Yang-Baxter algebra (see also [4]).
@@ 5.2. The set-theoretic algebras
(The decorated rack algebra.) Let ${\cal A}$ be the rack algebra (Definition 5.1). Let also $\sigma_{a},\ \tau_{b}:X\to X,$ and $\sigma_{a}$ be bijective for all $a\in X$. We say that the unital, associative algebra $\hat{\cal A}$ over $k,$ generated by indeterminates $1_{\hat{\cal A}},q_{a},q_{a}^{-1},h_{a},\in{\cal A}$ ($h_{a}=h_{b}\Leftrightarrow a=b$) and $w_{a},w^{-1}_{a}\in\hat{\cal A},$ $a\in X,$ and relations, $a,b\in X:$
$q_{a}^{-1}q_{a}=q_{a}q_{a}^{-1}=1_{\hat{\cal A}},\leavevmode\nobreak\ \leavevmode\nobreak\ q_{a}q_{b}=q_{b}q_{b\triangleright a},\leavevmode\nobreak\ \leavevmode\nobreak\ h_{a}h_{b}=\delta_{a,b}h_{a},\leavevmode\nobreak\ \leavevmode\nobreak\ q_{b}h_{b\triangleright a}=h_{a}q_{b},$
$w_{a}^{-1}w_{a}=w_{a}w_{a}^{-1}=1_{\hat{\cal A}},\leavevmode\nobreak\ \leavevmode\nobreak\ w_{a}w_{b}=w_{\sigma_{a}(b)}w_{\tau_{b}(a)}\leavevmode\nobreak\ \leavevmode\nobreak\ w_{a}h_{b}=h_{\sigma_{a}(b)}w_{a},\leavevmode\nobreak\ \leavevmode\nobreak\ w_{a}q_{b}=q_{\sigma_{a}(b)}w_{a}$ (5.17)
is a decorated rack algebra.
@@
Let $c=\sum_{a\in X}h_{a},$ then $c$ is a central element of the decorated rack algebra $\hat{\cal A}$.
@@
Let $\hat{\cal A}$ be the decorated rack algebra, then for all $a,b,c\in X,$
@@
It is worth pointing out the compatibility of the above proposition with Definition 2.14, which indicates the consistency of our construction.
@@
(The set-theoretic Yang-Baxter algebra.) Let ${\cal A}$ be the quandle algebra. Let also $\sigma_{a},\ \tau_{b}:X\to X,$ and $\sigma_{a}$ be bijective for all $a\in X.$ We say that the unital, associative algebra $\hat{\cal A}$ over $k,$ generated by indeterminates $1_{\hat{\cal A}},q_{a},q_{a}^{-1},h_{a},\in{\cal A}$ ($h_{a}=h_{b}\Leftrightarrow a=b$) and $w_{a},w^{-1}_{a}\in\hat{\cal A},$ $a\in X,$ and relations, (5.17) is a set-theoretic Yang-Baxter algebra.
@@
(Hopf algebra) Let $\hat{\cal A}$ be the set-theoretic Yang-Baxter algebra, ${\cal R}=\sum_{b\in X}h_{b}\otimes q_{b}$ be the rack universal ${\cal R}$-matrix, and $({\cal A},\Delta,\epsilon,S,{\cal R})$ be the quasi-triangular Hopf algebra of Theorem 5.12. Let also for all $a,b,x\in X,$
@@
Corollary 5.20.
Let $\hat{\cal A}$ be the decorated rack algebra, then the subalgebra consisting of the indeterminates $1_{\hat{\cal A}},\ q_{a}^{\pm 1}\ w_{a}^{\pm 1},$ and relations for all $a\in X$
$q_{a}^{-1}q_{a}=q_{a}q_{a}^{-1}=1_{\cal A},\leavevmode\nobreak\ \leavevmode\nobreak\ q_{a}q_{b}=q_{b}q_{b\triangleright a},$
$w_{a}^{-1}w_{a}=w_{a}w^{-1}_{a}=1_{\hat{\cal A}}\leavevmode\nobreak\ \leavevmode\nobreak\ w_{a}w_{b}=w_{\sigma_{a}(b)}w_{\tau_{b}(a)},\leavevmode\nobreak\ \leavevmode\nobreak\ w_{a}q_{b}=q_{\sigma_{a}(b)}w_{a}$ (5.19)
is a Hopf algebra with:
  1. (1)
    Co-product. $\Delta:{\cal A}\to{\cal A}\otimes{\cal A},$ $\leavevmode\nobreak\ \Delta(q_{a}^{\pm 1})=q_{a}^{\pm 1}\otimes q_{a}^{\pm 1},$ $\leavevmode\nobreak\ \Delta(w_{a}^{\pm 1})=w_{a}^{\pm 1}\otimes w_{a}^{\pm 1}.$
  2. (2)
    Co-unit. $\epsilon:{\cal A}\to k,$ $\leavevmode\nobreak\ \epsilon(q_{a}^{\pm 1})=1,$ $\leavevmode\nobreak\ \epsilon(w_{a}^{\pm 1})=1.$
  3. (3)
    Antipode. $S:{\cal A}\to{\cal A},$ $\leavevmode\nobreak\ S(q_{a}^{\pm 1})=q_{a}^{\mp 1}.$
Proof.
The proof is straightforward by Theorem 5.12 and Proposition 5.19. ∎
@@
Proposition 5.19 can be generalized for $(X,\bullet)$ being a magma , if for all $x\in X,$ $\exists g_{x}:X\to X,$ $a\mapsto g_{x}(a),$ being a bijection, such that for all $a,b,x\in X,$ there exists $e\in X,$ $\sigma_{e}(b)=b=g_{e}(b)$ and
@@
Note that in this case due to condition (5.20) we conclude that
$\Delta(w_{a})\Delta(h_{b})=\Delta(h_{g_{a}(b)})\Delta(w_{a}),$ (5.21)
that is in this case $\Delta:X\to X\otimes X$ is not an $\hat{\cal A}$ algebra homomorphism anymore.
We now introduce some handy notation that can be used in the following: let $i,j,k\in\{1,2,3\},$ then ${\cal F}_{jik}=\pi_{ij}\circ{\cal F}_{ijk}$ and ${\cal F}_{ikj}=\pi_{jk}\circ{\cal F}_{ijk},$ where $\pi$ is the flip map.
@@
(Drinfel’d twist [27]) Let ${\cal R}=\sum_{a\in X}h_{a}\otimes q_{a}\in{\cal A}\otimes{\cal A}$ be the rack universal ${\cal R}-$matrix. Let also $\hat{\cal A}$ be the decorated rack algebra and ${\cal F}\in\hat{\cal A}\otimes\hat{\cal A},$ such that ${\cal F}=\sum_{b\in X}h_{b}\otimes w_{b}^{-1},$ ${\cal R}_{ij}^{F}:={\cal F}_{ji}{\cal R}_{ij}{\cal F}_{ij}^{-1},$ $i,j\in\{1,2,3\}.$ We also define:
@@
    Indeed, this is proved by a direct computation and use of the decorated rack algebra. In fact, ${\cal F}_{123}=\sum_{a,b\in X}h_{a}\otimes h_{b}w_{a}^{-1}\otimes w_{b}^{-1}w_{a}^{-1}.$
@@
    (ii) By means of the relations of the decorated rack algebra $\hat{\cal A}$ we compute:
    ${\cal F}^{*}_{21,3}{\cal R}_{12}=\sum_{a,c\in X}h_{a}\otimes q_{a}h_{a\triangleright c}\otimes(w_{c}w_{\sigma^{-1}_{c}(a)})^{-1},\quad{\cal R}_{12}{\cal F}^{*}_{12,3}=\sum_{a,b\in X}h_{a}\otimes q_{a}h_{\sigma_{a}(b)}\otimes(w_{a}w_{b})^{-1}.$
    Due to the fact that $\leavevmode\nobreak\ b\triangleright a=\sigma_{b}(\tau_{\sigma^{-1}_{a}(b)}(a))$ and $w_{a}w_{b}=w_{\sigma_{a}(b)}w_{\tau_{b}(a)}$ we conclude that ${\cal F}^{*}_{21,3}{\cal R}_{12}={\cal R}_{12}{\cal F}^{*}_{12,3}$ and consequently (recall ${\cal F}_{213}={\cal F}_{21}{\cal F}^{*}_{21,3}$)
    ${\cal F}_{213}{\cal R}_{12}={\cal F}_{21}{\cal F}^{*}_{21,3}{\cal R}_{12}={\cal F}_{21}{\cal R}_{12}{\cal F}^{*}_{12,3}={\cal R}_{12}^{F}{\cal F}_{123}.\hfill\qed$
@@
[27] Let ${\cal F}$ be an admissible twist and ${\cal R}$ be a solution of the Yang-Baxter equation. Then ${\cal R}^{F}:={\cal F}^{(op)}{\cal R}{\cal F}^{-1}$ (${\cal F}^{(op)}=\pi\circ{\cal F},$ $\pi$ is the flip map) is also a solution of the Yang-Baxter equation.
@@
We note the direct correspondence of the above Proposition and Corollary with Theorem 2.15.
We are now going to examine the twisted ${\cal R}-$matrix as well as the twisted co-products (see also [21, 22, 23]). Henceforth, $\hat{\cal A}$ denotes the set-theoretic Yang-Baxter algebra.
@@
(Twisted universal ${\cal R}$-matrix) It is worth extracting the explicit expressions of the twisted universal ${\cal R}-$matrix and the twisted coproducts of the algebra. We recall the admissible twist ${\cal F}=\sum_{b\in X}h_{b}\otimes w_{b}^{-1}.$
@@
    Then, the twisted coproducts read as: $\leavevmode\nobreak\ \Delta_{F}(w_{a})=\sum_{b\in X}h_{\sigma_{a}(b)}w_{a}\otimes w_{\tau_{b}(a)},$
    $\Delta_{F}(h_{a})=\sum_{b\in X}h_{b}\otimes w_{b}^{-1}h_{c}w_{b}\Big{|}_{b\bullet c=a},$ $\leavevmode\nobreak\ \Delta_{F}(q_{a})=\sum_{b\in X}q_{a}h_{a\triangleright b}\otimes w_{b}^{-1}q_{a}w_{a\triangleright b},$